# A Law of Next-Token Prediction in Large Language Models

### v1

@@ 3 Discussion
The complex details of pre-trained LLMs have led many practitioners and researchers to anticipate that the internal workings of these models are inherently chaotic, lacking any consistent structures across different contexts. In this work, we challenge this view by introducing the law of equi-learning, which describes how contextualized token embeddings evolve from the first to the last layer. This law, which is both quantitative and precise, has been consistently observed across various architectures, including Transformer, Mamba, and RWKV. Its emergence provides crucial insights into the training and interpretation of LLMs, offering new perspectives that deepen our understanding of their internal mechanisms.
The significance of the equi-learning law lies in its potential to refine the development and application of LLMs. An open question is how the decay ratio $\rho$ depends on factors such as model depth and pre-training data. Understanding this dependence could lead to the development of more efficient LLMs by minimizing $\rho^{L-1}$, the overall decay ratio, in the equi-learning law. The law’s eventual emergence also suggests the possibility of setting different learning rates across layers to accelerate convergence to the equilibrium described by the law. Moreover, preserving the equi-learning law during model pruning and fine-tuning may yield practical benefits, potentially through the preservation of certain weights or the use of techniques like LoRA [18].
@@ Acknowledgments
We would like to thank Zhehang Du, Hangyu Lin, Cheng Shi, Peng Wang, and Yuan Yao for stimulating discussions. This research was supported in part by NSF grant DMS-2310679 and Wharton AI for Business.
Supplementary Materials
@@
Large language models (LLMs).
In this study, we focus on LLMs, where the objective is to predict the subsequent token in a sequence, constrained to attend solely to preceding tokens. Formally, given a sequence of tokens $x_{1}x_{2}\cdots x_{t}\in\mathcal{V}^{t}$, let $\mathbf{h}_{i,l}$ represent the contextualized token embeddings of the $i$-th token at the $l$-th layer, where $1\leq i\leq t,1\leq l\leq L$. The contextualized token embeddings from the top layer for the last token in the sequence, $\mathbf{h}_{t,L}$, is employed to predict the next token $x_{t+1}\in\mathcal{V}$, with $11\leq t\leq T-$ 1.
Prediction residual (PR).
@@
To assess the capability of contextualized token embeddings in predicting the next token, we calculate the fraction of variance unexplained (FVU) by a linear regression model predicting the next token. Notably, this metric is similar to the concurrent measure proposed by [36], specifically the Root Mean Squared Error (RMSE) of the optimal linear regressor based on the features in multilayer neural networks.
Layer normalization.
@@
In pre-layer normalization (pre-LN) models, default initialized layer normalization is applied to all layers except the last-layer token embeddings. This is because layer normalization is moved to the input of each sub-block, with an additional layer normalization added after the final self-attention block [31]. Given that different models utilize distinct forms of layer normalization—such as LayerNorm [2] in GPT-2 and RMSNorm [46] in Llama-1—we will apply the specific layer normalization technique used by each model to normalize its contextualized token embeddings. Throughout this paper, nearly all models are pre-LN models, with the exceptions of GPT-1, BERT, and RoBERTa.
Probing datasets.
@@
For our experiments, we consider eight distinct probing datasets: BookCorpus, C4, OpenWebText, Wikipedia, peS2o, Pile, Redpajama, and OSCAR. We sampled sentences based on their average length, extracting $3,000$ sentences from BookCorpus, $200$ from C4, and $100$ from each of the remaining datasets. For consistency, sentences were truncated to a maximum length of $512$ tokens across all datasets, except for BookCorpus. Unless otherwise specified, we will utilize the probing dataset that exhibits the strongest Pearson correlation in next-token prediction for each LLM.
@@
As depicted in Fig. 1, the law of equi-learning is observed in various open-source large language models. Please note that phi-1 was excluded from our analysis, as it is a LLM specifically designed for code. For each model, we evaluate the largest size that can be executed on a local machine equipped with two L40S GPUs, each possessing 48 GB of memory. The corresponding model sizes are as follows: 117M for GPT-1, 1.5B for GPT-2, 13B for Llama-1, 13B for Llama 2 and Llama 2-Chat, 8B for Llama 3 and Llama 3 Instruct, 7B for three versions of Mistral 7B and Mistral 7B-Instruct (v0.1, v0.2, v0.3), 1.3B for phi-1.5, 2.7B for phi-2, 14B for phi-3 (phi-3-medium) with varying context lengths (4K, 128K), 14B for RWKV and RWKV-Raven, and 2.8B for Mamba. The corresponding probing datasets used are as follows: BookCorpus for GPT-1, BookCorpus for GPT-2, OSCAR for Llama-1, peS2o for Llama 2 and Llama 2-Chat, BookCorpus for Llama 3 and Llama 3 Instruct, C4 for three versions of Mistral 7B and Mistral 7B-Instruct (v0.1, v0.2, v0.3), BookCorpus for phi-1.5, BookCorpus for phi-2, C4 for phi-3 (phi-3-medium) with varying context lengths (4K, 128K), C4 for RWKV and RWKV-Raven, and BookCorpus for Mamba. Notably, reinforcement learning from human feedback (RLHF) [27] does not significantly impact the law, as demonstrated by Llama-2-13B, Llama-3-8B, Mistral-7B-v0.1, Mistral-7B-v0.2, Mistral-7B-v0.3, and RWKV, along with their fine-tuned versions, as illustrated in Fig. 1.
@@
Fig. S1: Intermediate-layer contextualized token embeddings for they</w> (red) and them</w> (blue) plotted on the plane of the first two principal components.
@@
Fig. S2: Intermediate-layer contextualized token embeddings for have</w> (red) and had</w> (blue) plotted on the plane of the first two principal components.
@@
Fig. S3: Intermediate-layer contextualized token embeddings for are</w> (red) and is</w> (blue) plotted on the plane of the first two principal components.
@@
Fig. S4: The law of equi-learning emerges when GPT-1 is evaluated across different domains, including medicine (MedRAG-Textbooks), law (LegalBench), and politics (US-Congressional-Speeches), with corresponding Pearson correlation coefficients of $-0.990$, $-0.995$, and $-0.998$, respectively.
@@
Fig. S5: Intermediate-layer contextualized token embeddings for law</w> (red) and policy</w> (blue) plotted on the plane of the first two principal components.
@@
Fig. S6: Intermediate-layer contextualized token embeddings for president</w> (red) and country</w> (blue) plotted on the plane of the first two principal components.
Visualization.
@@
Figs. S1, S2, and S3 depict the visualization of contextualized token embeddings across various layers of the GPT-1 model, using $3,000$ sentences from the BookCorpus (consistent with the setting in Fig.1). We examine three token pairs: they</w> vs. them</w>, have</w> vs. had</w>, and are</w> vs. is</w>. Principal component analysis (PCA) is applied to project these contextualized token embeddings onto a two-dimensional plane. The results show a distinct and progressive separation of contextualized token embeddings within each pair as one moves from the lower to the upper layers of the model. Moreover, as illustrated in Fig. S4, the law of equi-learning manifests when evaluating the GPT-1 model across distinct domains. Specifically, we analyzed $200$ sentences from MedRAG-Textbooks [45] for the medicine domain, $1000$ sentences from LegalBench [14] for the law domain, and $500$ sentences from US-Congressional-Speeches [^2] See more in https://huggingface.co/datasets/Eugleo/us-congressional-speeches. for the politics domain. Consistent with other probing datasets, the number of sampled sentences was determined by their average length, and sentences were truncated to a maximum length of $512$ tokens. For each domain, we visualized the embeddings of token sets across various layers of the GPT-1 model—patients</w>, cells</w>, and disorder</w> for medicine; law</w> and policy</w> for law; and president</w> and country</w> for politics. Similarly, PCA was employed to project these contextualized token embeddings onto a two-dimensional plane. As depicted in Fig. 2, Fig. S5, and Fig. S6, the results reveal a distinct and progressively increasing separation of contextualized token embeddings within each token set as the model transitions from lower to upper layers.
Fig. S7: An enlarged version of Pythia-1B at initialization (Step=0).
Training dynamics.
@@
As illustrated in Figs. 3, 4, and 5, we examine the evolution of the observed law throughout the training process. Specifically, Fig. 3 depicts the PR of contextualiezed token embeddings at each layer of Pythia-1B at various training steps (an enlarged version at initialization (Step=0) is in Fig. S7), using BookCorpus as the probing dataset. Additionally, Figs. 4 and 5 present the PR of contextualized token embeddings at each layer of pre-trained 2.8B GPT-2 models, as released by [26], across different training epochs and data repetitions, with OpenWebText serving as the probing dataset.
Model scaling.
@@
As illustrated in Fig. 6, the contextualized token embeddings are compared across different model sizes within the same model series. We analyze three distinct model series—GPT-2, RWKV-Raven, and Mamba—each with at least four different model sizes. The largest version of each series is depicted in Fig. 1.
Pre-training task.
@@
As illustrated in Fig. 7, different probing tasks are employed to analyze the contextualized token embeddings of BERT, RoBERTa, and T5. Notably, the probing datasets used are peS2o for BERT and RoBERTa, and C4 for T5. Since masked language modeling (MLM) and span corruption (SC) mask or corrupt only 15% of tokens, leading to a corresponding reduction to 15% of the effective examples compared to NTP, we multiply the size of the probing datasets by 7 for MLM and SC.
Information flow.
As illustrated in Fig. 8, the contextualized token embeddings of the current token ($x_{t}$) at each layer are leveraged to predict various tokens in the sequence, spanning previous token ($x_{t-1}$) to next next token ($x_{t+2}$), including the default next token ($x_{t+1}$). For clarity, we present four models selected from those shown in Fig. 1: GPT2-XL, Llama-3-8B-Instruct, RWKV-Raven-14B, and Mamba-2.8B.
@@
Fig. S8: With the measure of separation fuzziness, the law of equi-learning is not clear though the decreasing trend is still observed.
@@
Fig. S9: When the vocabulary is shuffled, the law of equi-learning is not very clear though the decreasing trend is still observed.
Measure analysis.
Figure S8 illustrates that the widely used metric of separation fuzziness, commonly applied to assess deep learning features in classification tasks [28, 9, 16], is inadequate for the emergence of the law of equi-learning in LLMs. This inadequacy may stem from the larger token vocabulary size compared to the embedding dimension and the presence of very similar or even identical contexts followed by different tokens in natural language data [44]. For simplicity, we selected four models from Fig. 1: GPT-2 XL, Llama-3-8B-Instruct, RWKV-Raven-14B, and Mamba-2.8B, and utilized separation fuzziness to evaluate the quality of contextualized token embeddings. Furthermore, as demonstrated in Fig. S9, the law of equi-learning becomes obscured when the vocabulary is shuffled, resulting in differing token index orders. This observation suggests that the widely adopted byte pair encoding (BPE) algorithm [35] in tokenizers can produce a meaningful token index order, which is critical to the emergence of the law of equi-learning. Further investigation is needed and is deferred to future work.
@@
Fig. S10: For pre-LN models, omitting the additional initialized layer normalization (no LN) for contextualized token embeddings leads to a noisier manifestation of the law of equi-learning compared to the default approach with initialized layer normalization. This normalization effect can also be achieved through standardization, rather than relying on initialized layer normalization.
Layer normalization analysis.
As illustrated in Fig. S10, layer normalization plays a critical role in pre-LN models, as the absence of additional layer normalization makes the law noisier. Notably, this normalization effect can also be attained through standardization, specifically $z=\frac{x-\mu}{\sigma}$ across the embedding dimension. For simplicity, we selected four models from Fig. 1—GPT-2 XL, Llama-3-8B-Instruct, RWKV-Raven-14B, and Mamba-2.8B—and analyzed the impact of layer normalization. This finding aligns with observations related to the equi-separation law in multilayer neural networks[16], where batch normalization is crucial for its emergence. It is important to note that batch normalization does not impact the PR of contextualized token embeddings in our case.
@@
Fig. S11: With unsuitable probing dataset (e.g., RedPajama here), the law of equi-learning is not clear, though a descending trend is still observed.
Probing data analysis.
As demonstrated in Fig. S11, the observed law’s clarity diminishes when an inappropriate probing dataset is selected, although a general descending trend remains evident. For simplicity, we selected four models from Fig. 1—GPT-2 XL, Llama-3-8B-Instruct, RWKV-Raven-14B, and Mamba-2.8B—and presented the PR of their token embeddings using the RedPajama dataset as the probing dataset. Moreover, the optimal probing dataset among the eight evaluated shifts from OSCAR and peS2o to BookCorpus as models progress from Llama-1 and Llama 2 (including its chat variant) to Llama 3 (and its instruct variant), suggesting that higher-quality pre-training data may necessitate higher-quality probing data to facilitate the emergence of the law of equi-learning. These findings underscore the critical importance of selecting appropriate probing data for the emergence of the law of equi-learning.

### v3

@@ 4 Discussion
Despite the extensive research on structures within pre-trained LLMs, many have found it challenging to identify precise, quantitative laws governing their internal dynamics. In this work, we challenge this view by introducing the law of equi-learning, which describes how contextualized token embeddings evolve from the first to the last layer. This law, which is both quantitative and precise, has been consistently observed across various architectures, including Transformer, Mamba, and RWKV. Its emergence provides crucial insights into the training and interpretation of LLMs, offering new perspectives that deepen our understanding of their internal mechanisms.
The significance of the equi-learning law lies in its potential to refine the development and application of LLMs. An open question is how the decay ratio $\rho$ depends on factors such as model depth and pre-training data. Understanding this dependence could lead to the development of more efficient LLMs by minimizing $\rho^{L-1}$, the overall decay ratio, in the equi-learning law. The emergence of this law specifically under the PR metric defined in (2), rather than alternative metrics (see Fig. S8 in the Supplementary Materials), warrants further investigation. We hypothesize that this specificity may arise from the PR metric’s incorporation of token indices derived from byte pair encoding [37], which might capture structural information beyond simple classification metrics. Nevertheless, we believe that the law could be extended beyond the PR metric, though this remains an avenue for future research. The law’s eventual emergence also suggests the possibility of setting different learning rates across layers to accelerate convergence to the equilibrium described by the law. Moreover, preserving the equi-learning law during model pruning and fine-tuning may yield practical benefits, potentially through the preservation of certain weights or the use of techniques like LoRA [20].
A central open question is to uncover the mechanism underlying the equi-learning law. Related phenomena have been analytically derived in deep linear networks (DLNs) with linearly separable data [46], but the strong assumptions—such as the absence of nonlinearity and simplified data—limit their relevance to real-world LLMs. A spring-block analogy has been proposed to illustrate a similar law in deep neural networks (DNNs) [38]. However, this framework remains primarily heuristic, as it does not establish a concrete correspondence between fundamental physical elements—such as elastic potential energy and friction—and specific components within DNNs. Furthermore, it does not address the substantial architectural and functional disparities between DNNs and modern LLMs. A formal derivation of the equi-learning law in LLMs will likely require a deeper understanding of model dynamics and the structure of data, aligned with recent geometric approaches to analyzing DNNs and LLMs [30, 48, 4, 50], which we leave for future work.
@@ Acknowledgments
We would like to thank Zhehang Du, Hangyu Lin, Cheng Shi, Peng Wang, and Yuan Yao for stimulating discussions. We are grateful to the two anonymous reviewers whose comments helped improve the presentation of this paper. This research was supported in part by NSF grant DMS-2310679 and Wharton AI for Business.
Supplementary Text
@@
LLMs. In this study, we focus on autoregressive LLMs, where the objective is to predict the subsequent token in a sequence, constrained to attend solely to preceding tokens. Formally, the model takes as input a sequence of discrete tokens $x_{1},x_{2},\dots,x_{t}\in\mathcal{V}^{t}$, where $\mathcal{V}$ denotes the vocabulary specific to the model. These tokens are obtained via a tokenizer applied to raw text; for instance, the sentence “We love Physics.” is tokenized by the Llama 3 tokenizer into four tokens: “We”, “ love”, “ Physics”, and “.” Vocabulary sizes vary across models (e.g., 32K for Llama 2 and 128K for Llama 3). Each token $x_{i}$ is mapped to an initial embedding vector $\mathbf{h}_{i,0}$ via a learned embedding table. These embeddings are then propagated through a stack of model layers—typically transformer layers—resulting in a set of contextualized token embeddings $\mathbf{h}_{i,\ell}$ at each layer $1\leq\ell\leq L$. Due to the autoregressive nature of the model, the contextualized token embeddings $\mathbf{h}_{i,\ell}$ is computed using only information from position $i$ and all positions preceding it, specifically from the previous layer’s outputs $\{\mathbf{h}_{j,\ell-1}\mid 1\leq j\leq i\}$. This ensures that the model does not access future tokens. The last-layer embedding of the current token, $\mathbf{h}_{t,L}$ (also denoted $\mathbf{h}_{t,\text{last}}$), is used to predict the next token $x_{t+1}$ by projecting it into vocabulary space and applying a softmax to produce a probability distribution over the vocabulary. This autoregressive decoding procedure is iteratively applied at each position $1\leq t\leq T-1$, where $T$ denotes the length of the input token sequence.
@@
Prediction residual (PR). To assess the capability of contextualized token embeddings in predicting the next token, we calculate the fraction of variance unexplained (FVU) by a linear regression model predicting the next token. Notably, this metric is similar to the concurrent measure proposed by [38], specifically the Root Mean Squared Error (RMSE) of the optimal linear regressor based on the features in MLPs. In their study, RMSE was shown to reproduce a similar noise–nonlinearity phase diagram in MLP training under regression, extending phenomena originally observed in classification to the regression setting. This finding supports the potential of PR as a meaningful measure of representation quality.
@@
Layer normalization. In pre-layer normalization (pre-LN) models, default initialized layer normalization is applied to all layers except the last-layer token embeddings. This is because layer normalization is moved to the input of each sub-block, with an additional layer normalization added after the final self-attention block [33]. Given that different models utilize distinct forms of layer normalization—such as LayerNorm [2] in GPT-2 and RMSNorm [51] in Llama-1—we will apply the specific layer normalization technique used by each model to normalize its contextualized token embeddings. Throughout this paper, nearly all models are pre-LN models, with the exceptions of GPT-1, BERT, and RoBERTa.
@@
Probing datasets. For our experiments, we consider eight distinct probing datasets: BookCorpus, C4, OpenWebText, Wikipedia, peS2o, Pile, Redpajama, and OSCAR. We sampled sentences based on their average length, extracting $3,000$ sentences from BookCorpus, $200$ from C4, and $100$ from each of the remaining datasets. For consistency, sentences were truncated to a maximum length of $512$ tokens across all datasets, except for BookCorpus. Unless otherwise specified, we will utilize the probing dataset that exhibits the strongest Pearson correlation in next-token prediction for each LLM.
@@
Large language models. As depicted in Fig. 1, the law of equi-learning is observed in various open-source large language models. Please note that phi-1 was excluded from our analysis, as it is a LLM specifically designed for code. For each model, we evaluate the largest size that can be executed on a local machine equipped with two L40S GPUs, each possessing 48 GB of memory. The corresponding model sizes are as follows: 117M for GPT-1, 1.5B for GPT-2, 13B for Llama-1, 13B for Llama 2 and Llama 2-Chat, 8B for Llama 3 and Llama 3 Instruct, 7B for three versions of Mistral 7B and Mistral 7B-Instruct (v0.1, v0.2, v0.3), 1.3B for phi-1.5, 2.7B for phi-2, 14B for phi-3 (phi-3-medium) with varying context lengths (4K, 128K), 14B for RWKV and RWKV-Raven, and 2.8B for Mamba. The corresponding probing datasets used are as follows: BookCorpus for GPT-1, BookCorpus for GPT-2, peS2o for Llama-1, peS2o for Llama 2 and Llama 2-Chat, BookCorpus for Llama 3 and Llama 3 Instruct, C4 for three versions of Mistral 7B and Mistral 7B-Instruct (v0.1, v0.2, v0.3), BookCorpus for phi-1.5, BookCorpus for phi-2, C4 for phi-3 (phi-3-medium) with varying context lengths (4K, 128K), C4 for RWKV and RWKV-Raven, and BookCorpus for Mamba. Notably, reinforcement learning from human feedback (RLHF) [29] does not significantly impact the law, as demonstrated by Llama-2-13B, Llama-3-8B, Mistral-7B-v0.1, Mistral-7B-v0.2, Mistral-7B-v0.3, and RWKV, along with their fine-tuned versions, as illustrated in Fig. 1.
@@
Fig. S1: Intermediate-layer contextualized token embeddings for they</w> (red) and them</w> (blue) plotted on the plane of the first two principal components. The x-axis and y-axis represent the first and second principal components, respectively.
@@
Fig. S2: Intermediate-layer contextualized token embeddings for have</w> (red) and had</w> (blue) plotted on the plane of the first two principal components. The x-axis and y-axis represent the first and second principal components, respectively.
@@
Fig. S3: Intermediate-layer contextualized token embeddings for are</w> (red) and is</w> (blue) plotted on the plane of the first two principal components. The x-axis and y-axis represent the first and second principal components, respectively.
@@
Fig. S4: The law of equi-learning emerges when GPT-1 is evaluated across different domains, including medicine (MedRAG-Textbooks), law (LegalBench), and politics (US-Congressional-Speeches), with corresponding Pearson correlation coefficients of 0.990-0.990, 0.995-0.995, and 0.998-0.998, respectively. The x-axis denotes the layer index, while the y-axis (log scale) shows the prediction residual (PR) as defined in Eq. 2.
@@
Fig. S5: Intermediate-layer contextualized token embeddings for law</w> (red) and policy</w> (blue) plotted on the plane of the first two principal components. The x-axis and y-axis represent the first and second principal components, respectively.
@@
Fig. S6: Intermediate-layer contextualized token embeddings for president</w> (red) and country</w> (blue) plotted on the plane of the first two principal components. The x-axis and y-axis represent the first and second principal components, respectively.
@@
Visualization. Figs. S1, S2, and S3 depict the visualization of contextualized token embeddings across various layers of the GPT-1 model, using $3,000$ sentences from the BookCorpus (consistent with the setting in Fig.1). We examine three token pairs: they</w> vs. them</w>, have</w> vs. had</w>, and are</w> vs. is</w>. Principal component analysis (PCA) is applied to project these contextualized token embeddings onto a two-dimensional plane. The results show a distinct and progressive separation of contextualized token embeddings within each pair as one moves from the lower to the upper layers of the model. Moreover, as illustrated in Fig. S4, the law of equi-learning manifests when evaluating the GPT-1 model across distinct domains. Specifically, we analyzed $200$ sentences from MedRAG-Textbooks [49] for the medicine domain, $1000$ sentences from LegalBench [15] for the law domain, and $500$ sentences from US-Congressional-Speeches [^3] See more in https://huggingface.co/datasets/Eugleo/us-congressional-speeches. for the politics domain. Consistent with other probing datasets, the number of sampled sentences was determined by their average length, and sentences were truncated to a maximum length of $512$ tokens. For each domain, we visualized the embeddings of token sets across various layers of the GPT-1 model—patients</w>, cells</w>, and disorder</w> for medicine; law</w> and policy</w> for law; and president</w> and country</w> for politics. Similarly, PCA was employed to project these contextualized token embeddings onto a two-dimensional plane. As depicted in Fig. 2, Fig. S5, and Fig. S6, the results reveal a distinct and progressively increasing separation of contextualized token embeddings within each token set as the model transitions from lower to upper layers.
Fig. S7: An enlarged version of Pythia-1B at initialization (Step=0). The x-axis denotes the layer index, while the y-axis (log scale) shows the prediction residual (PR) as defined in Eq. 2.
@@
Training dynamics. As illustrated in Figs. 3, 4, and 5, we examine the evolution of the observed law throughout the training process. Specifically, Fig. 3 depicts the PR of contextualiezed token embeddings at each layer of Pythia-1B at various training steps (an enlarged version at initialization (Step=0) is in Fig. S7), using BookCorpus as the probing dataset. Additionally, Figs. 4 and 5 present the PR of contextualized token embeddings at each layer of pre-trained 2.8B GPT-2 models, as released by [28], across different training epochs and data repetitions, with OpenWebText serving as the probing dataset.
@@
Model scaling. As illustrated in Fig. 6, the contextualized token embeddings are compared across different model sizes within the same model series. We analyze three distinct model series—GPT-2, RWKV-Raven, and Mamba—each with at least four different model sizes. The largest version of each series is depicted in Fig. 1.
@@
Pre-training task. As illustrated in Fig. 7, different probing tasks are employed to analyze the contextualized token embeddings of BERT, RoBERTa, and T5. Notably, the probing datasets used are peS2o for BERT and RoBERTa, and C4 for T5. Since masked language modeling (MLM) and span corruption (SC) mask or corrupt only 15% of tokens, leading to a corresponding reduction to 15% of the effective examples compared to NTP, we multiply the size of the probing datasets by 7 for MLM and SC.
Information flow. As illustrated in Fig. 8, the contextualized token embeddings of the current token ($x_{t}$) at each layer are leveraged to predict various tokens in the sequence, spanning previous token ($x_{t-1}$) to next next token ($x_{t+2}$), including the default next token ($x_{t+1}$). For clarity, we present four models selected from those shown in Fig. 1: GPT2-XL, Llama-3-8B-Instruct, RWKV-Raven-14B, and Mamba-2.8B. For simplicity, we present an explicit formulation for predicting the previous token, $x_{t-1}$; analogous formulas for predicting other tokens can be obtained by appropriately adjusting the prediction target in $\text{PR}^{\text{prev}}$ and $\text{PR}^{\text{prev}}_{\ell}$. Specifically, the model uses the last-layer embedding of the current token, denoted as $\mathbf{h}_{t,\text{last}}$, to predict the previous token in the sequence. Let $x^{s}_{t-1}$ represent the immediately preceding token that the LLM aims to predict based on the first $t$ tokens, $x^{s}_{1},x^{s}_{2},\ldots,x^{s}_{t}$, through the last-layer embedding of the current token $\mathbf{h}_{t,\text{last}}^{s}$, for an index $1\leq s\leq S$ over all training corpus. This process forms a dataset $\mathcal{D}_{\text{prev}}:=\{(\mathbf{h}_{t,\text{last}}^{s},x^{s}_{t-1})\mid 1\leq s\leq S\}$. To assess the capability of the LLM’s current token embeddings in predicting the previous token, we evaluate how well a linear regression model fits on the dataset $\mathcal{D}_{\text{prev}}$. For this purpose, we identify $x$ with its index in the token vocabulary. Let $\hat{x}_{\text{prev}}=\mathbf{w}\cdot\mathbf{h}+b$ denote the least-squares fit on $\mathcal{D}_{\text{prev}}$. This suggests using the following metric to quantify the LLM’s previous-token prediction capability:
$\text{PR}^{\text{prev}}:=\frac{\sum(x_{\text{prev}}-\hat{x}_{\text{prev}})^{2}}{\sum(x_{\text{prev}}-\bar{x}_{\text{prev}})^{2}},$ (2)
where the sum is over all $x_{\text{prev}}=x^{s}_{t-1}$, $\hat{x}_{\text{prev}}=\mathbf{w}\cdot\mathbf{h}_{t,\text{last}}^{s}+b$, and $\bar{x}_{\text{prev}}$ represents the mean of all $x^{s}_{t-1}$. To investigate how the predictive power of an LLM with depth $L$ evolves across its layers, we calculate the $\text{PR}^{\text{prev}}$ for the previous-token prediction task at each intermediate layer. Let $\text{PR}^{\text{prev}}_{l}$ denote this value for the $l$-th layer, where $1\leq l\leq L$. Specifically, instead of using the last-layer embedding $\mathbf{h}_{t,\text{last}}\equiv\mathbf{h}_{t,L}$, we use the embedding of the current token at layer $\ell$, denoted $\mathbf{h}_{t,\ell}$, to predict the previous token when computing $\text{PR}^{\text{prev}}_{\ell}$.
@@
Fig. S8: With the measure of separation fuzziness, the law of equi-learning is not clear though the decreasing trend is still observed. The x-axis denotes the layer index, while the y-axis (log scale) shows the separation fuzziness.
@@
Fig. S9: When the vocabulary is shuffled, the law of equi-learning is not very clear though the decreasing trend is still observed. The x-axis denotes the layer index, while the y-axis (log scale) shows the prediction residual (PR) as defined in Eq. 2.
Measure analysis. Fig. S8 illustrates that the widely used metric of separation fuzziness, commonly applied to assess deep learning features in classification tasks [30, 10, 17], is inadequate for the emergence of the law of equi-learning in LLMs. This inadequacy may stem from the larger token vocabulary size compared to the embedding dimension and the presence of very similar or even identical contexts followed by different tokens in natural language data [48]. For simplicity, we selected four models from Fig. 1: GPT-2 XL, Llama-3-8B-Instruct, RWKV-Raven-14B, and Mamba-2.8B, and utilized separation fuzziness to evaluate the quality of contextualized token embeddings. Furthermore, as demonstrated in Fig. S9, the law of equi-learning becomes obscured when the vocabulary is shuffled, resulting in differing token index orders. This observation suggests that the widely adopted byte pair encoding algorithm [37] in tokenizers can produce a meaningful token index order, which is critical to the emergence of the law of equi-learning. Further investigation is needed and is deferred to future work.
@@
Fig. S10: For pre-LN models, omitting the additional initialized layer normalization (no LN) for contextualized token embeddings leads to a noisier manifestation of the law of equi-learning compared to the default approach with initialized layer normalization. This normalization effect can also be achieved through standardization, rather than relying on initialized layer normalization. The x-axis denotes the layer index, while the y-axis (log scale) shows the prediction residual (PR) as defined in Eq. 2.
Layer normalization analysis. As illustrated in Fig. S10, layer normalization plays a critical role in pre-LN models, as the absence of additional layer normalization makes the law noisier. Notably, this normalization effect can also be attained through standardization, specifically $z=\frac{x-\mu}{\sigma}$ across the embedding dimension. For simplicity, we selected four models from Fig. 1—GPT-2 XL, Llama-3-8B-Instruct, RWKV-Raven-14B, and Mamba-2.8B—and analyzed the impact of layer normalization. This finding aligns with observations related to the equi-separation law in MLPs [17], where batch normalization is crucial for its emergence. It is important to note that batch normalization does not impact the PR of contextualized token embeddings in our case.
[…]
GPT-1 -0.997 -0.951 -0.959 -0.943 -0.972 -0.917 -0.884 -0.969
GPT-2 XL -0.994 -0.960 -0.982 -0.962 -0.963 -0.965 -0.962 -0.957
Llama-1-13B -0.956 -0.993 -0.990 -0.986 -0.994 -0.992 -0.993 -0.993
Llama-2-13B -0.913 -0.985 -0.977 -0.966 -0.988 -0.973 -0.983 -0.985
Llama-2-13B-Chat -0.879 -0.984 -0.967 -0.962 -0.983 -0.964 -0.978 -0.980
Llama-3-8B -0.993 -0.981 -0.977 -0.979 -0.938 -0.941 -0.948 -0.969
Llama-3-8B-Instruct -0.992 -0.981 -0.974 -0.979 -0.940 -0.940 -0.948 -0.967
Mistral-7B-v0.1 -0.874 -0.988 -0.950 -0.968 -0.956 -0.941 -0.940 -0.979
Mistral-7B-Instruct-v0.1 -0.850 -0.991 -0.958 -0.971 -0.961 -0.939 -0.948 -0.985
Mistral-7B-v0.2 -0.863 -0.989 -0.952 -0.967 -0.953 -0.936 -0.941 -0.979
Mistral-7B-Instruct-v0.2 -0.874 -0.988 -0.952 -0.967 -0.958 -0.931 -0.936 -0.979
Mistral-7B-v0.3 -0.863 -0.989 -0.952 -0.967 -0.953 -0.936 -0.941 -0.979
Mistral-7B-Instruct-v0.3 -0.863 -0.988 -0.951 -0.966 -0.951 -0.932 -0.938 -0.979
phi-1.5 -0.994 -0.967 -0.986 -0.976 -0.974 -0.958 -0.967 -0.971
phi-2 -0.993 -0.984 -0.978 -0.983 -0.890 -0.954 -0.930 -0.987
phi-3-medium-4k-instruct -0.894 -0.992 -0.961 -0.955 -0.977 -0.972 -0.975 -0.991
phi-3-medium-128k-instruct -0.902 -0.992 -0.962 -0.959 -0.975 -0.972 -0.974 -0.991
RWKV-14B -0.984 -0.991 -0.983 -0.987 -0.967 -0.984 -0.952 -0.983
RWKV-Raven-14B -0.984 -0.991 -0.982 -0.984 -0.964 -0.984 -0.952 -0.981
[…]
Table S1: Pearson correlation coefficients between the logarithm of PR values (Eq. 2) and layer indices for all LLMs in Fig. 1, evaluated across eight probing datasets. For each LLM, the probing dataset with the strongest Pearson correlation is highlighted in bold.
@@
Fig. S11: With unsuitable probing dataset (e.g., RedPajama here), the law of equi-learning is not clear, though a descending trend is still observed. The x-axis denotes the layer index, while the y-axis (log scale) shows the prediction residual (PR) as defined in Eq. 2.
Probing data analysis. Table S1 highlights that the probing datasets exhibiting the strongest Pearson correlations for various LLMs in Fig.1 are BookCorpus, C4, and peS2o among the eight evaluated datasets. This trend may be attributed to the high quality of these datasets and their resemblance to the pre-training data used for these models. Notably, the optimal probing dataset among the eight evaluated transitions from peS2o to BookCorpus as models progress from Llama-1 and Llama 2 (including its chat variant) to Llama 3 (and its instruct variant). This shift suggests that higher-quality pre-training data may require correspondingly higher-quality probing datasets to facilitate the emergence of the law of equi-learning, likely reflecting the impact of carefully designed pre-processing and curation pipelines used to enhance the pre-training data quality in Llama 3 [9]. Furthermore, as demonstrated in Fig. S11, the observed law’s clarity diminishes when an inappropriate probing dataset is selected, although a general descending trend remains evident. For simplicity, we selected four models from Fig. 1—GPT-2 XL, Llama-3-8B-Instruct, RWKV-Raven-14B, and Mamba-2.8B—and presented the PR of their token embeddings using the RedPajama dataset as the probing dataset. These findings underscore the critical importance of selecting appropriate probing data for the emergence of the law of equi-learning.